1. Presented Below are selected transactions for ABC Company during  June 2021. ABC uses the perpetual inventory system.


June  4 Purchase merchandise from Mr. X at a cost of taka 45000, terms FOB shipping point, 3/10, n/30.

June  5 Paid freight charges of taka 2000 on June 4 merchandise purchase

June  7 Returned damaged goods costing taka 7000 received from Mr. X 

June 13 Sold merchandise to Mr. Y costing taka 15000 on accounts for 21000, terms FOB destination point 1/10, n/30.

June 14 Paid Mr. X due related to June 4 transactions.

June   15 Paid freight on June 13 sale taka 2500.

June 16 Made refunds to Mr. Y for defective Merchandise taka 2100

June 21 Received balance due from Mr. Y.

    

Requirement: Journalize the transactions using net basis.


1
Expert's answer
2021-09-14T06:05:47-0400

Computation of Accounts Payable (Mr. X) because payment has been made within 10 days to got the discount of 3%

Net amount paid to Mr X"= Purchases – Return \times (\frac{100 – Discount \space Rate}{100})\\

 = (45,000 – 7,000) \times (\frac{100 – 3}{100})\\

 = 38,000 \times{ 97}{100}\\

 = 36,860"

 

Computation of Accounts Receivables (Mr. Y) because payment has been received within 10 days to discount has been given 1%

"Net amount received from Mr Y \\= (Sales – Return) \times (100 – \frac{Discount\space Rate}{100})\\

 = (21,000 – 2,100) \times (\frac{100 – 1}{100})\\

= 18,900 \times \frac{ 99}{100}\\
 = 187,11"

 

Journal Entries



continuation of the table